OpenStudy (anonymous):

Simplify: y^2=2y+3y^2

OpenStudy (anonymous):

^ is to the power of by the way

OpenStudy (babalooo):

collect like terms and then simplify

OpenStudy (anonymous):

Is 2y a like term of y squared and 3y squared?

OpenStudy (anonymous):

Would it be 3y^3+2y or 4y^3+2y or something different?

OpenStudy (babalooo):

okay. sorry for the late reply, i was helping someone else. remember this. the coefficient in front like 2y, 3y simply means y+y =2y y+y+y=3y the the power like y^2 and y^3 means y*y=y^2 y*y*y=y^3 2y^2 means anything like 2*y*y or 2*y^2 or y^2+y^2

OpenStudy (babalooo):

only terms with variables to the same power can be added or subtracted. 2y+y= 3y 2y^2 +3y^2 +5y= 5y^2+5y